You are an option speculator. The current market price of a diamond is 1,000 USD and you expect it to rise in 2 years to 1,500 USD. On the market you can purchase CALL and PUT ike price of 1,600 USD and pre Which option should you buy based on your predictions? What would be your actual result if the market price of a diamond at delivery date was 1,530 USD? Result: P/L = 20 USD
Answer #1

if you are a speculator and expect the price to rise, then to safeguard from fall in price, you will buy put option.

Answer 1 : Put option

If market price of diamond on delivery date was 1530, then there will a payoff of 1600-1530 = 70 and premium paid is 50

so net profit from buying put option = 70 - 50 = 20

Answer 2 : Profit of $20 (Any doubts, feel free to ask. THUMBS UP PLEASE)